Find the sixth term of the sequence
2, 6, 12, 20, 30,...

Found 2 solutions by AnlytcPhil, josgarithmetic:
Answer by AnlytcPhil(1806)

2, 6, 12, 20, 30,...

 2 = 1×2
 6 = 2×3
12 = 3×4
20 = 4×5
30 = 5×6

So the next term is 6×7 = 42.

The nth term is 

an = n(n+1)

Edwin


Answer by josgarithmetic(39618)

A clear relationship is happening over the differences between successive terms.

These are those differences, as a list: 4, 6, 8, 10,...
The next difference should be 12. This will be the term 30+12=42.
The next difference should be 14. This will give the term 42+14=56.